How many grams of salt must be added to 80 grams of a 22% salt solution to obtain a 40% salt solution? show work.
Answer by checkley77(12844)

.22*80+X=.40(80+X)
17.6+X=32+.40X
X-.4X=32-17.6
.6X=14.4
X=14.4/.6
X=24 GRAMS OF SALT ARE ADDED.
PROOF:
.22*80+24=.40(80+24)
17.6+24=.40*104
41.6=41.6